A ferryboat, whose speed in still water is 3.20m/s, must cross a 340m wide river and arrive at a point 130m upstream from where it starts. To do so, the pilot must point the boat upstream mat an angle of 47 ̊. What is the speed of the river’s current?
1
Expert's answer
2020-12-17T07:37:54-0500

"v_y=3.2\cdot\cos47°=2.18(m/s)"


The time of traveling


"t=\frac{340}{v_y}=\frac{340}{2.18}=156(s)"


"v_x=3.2\cdot\sin47°=2.34(m/s)"


"(v_x-u)t=130\to u=v_x-130/t=2.34-130/156=1.5(m/s)". Answer